### v1

# T cell receptor binding prediction: A machine learning revolution

Abstract
Recent advancements in immune sequencing and experimental techniques are generating extensive T cell receptor (TCR) repertoire data, enabling the development of models to predict TCR binding specificity. Despite the computational challenges due to the vast diversity of TCRs and epitopes, significant progress has been made. This paper discusses the evolution of the computational models developed for this task, with a focus on machine learning efforts, including the early unsupervised clustering approaches, supervised models, and the more recent applications of Protein Language Models (PLMs). We critically assess the most prominent models in each category, and discuss recurrent challenges, such as the lack of generalization to new epitopes, dataset biases, and biases in the validation design of the models.
Furthermore, our paper discusses the transformative role of transformer-based protein models in bioinformatics. These models, pretrained on extensive collections of unlabeled protein sequences, can convert amino acid sequences into vectorized embeddings that capture important biological properties. We discuss recent attempts to leverage PLMs to deliver very competitive performances in TCR-related tasks. Finally, we address the pressing need for improved interpretability in these often opaque models, proposing strategies to amplify their impact in the field.
Keyword: Machine Learning; T cell Receptor; Specificity Prediction; Protein Language Models; Interpretability.
1 Background
T cells are an essential component of the adaptive immune system, due to their ability to orchestrate targeted, effective immune responses through cell-based and cytokine-release mechanisms. While T cell functions are diverse, their activation, differentiation, proliferation, and function are all governed by their T cell receptors (TCR), which enable them to recognize non-self antigens arising from infectious agents or diseased cells [1].
To face a diverse and ever-evolving array of antigens, the immune system has evolved the capability to generate a huge array of distinct TCRs. This diversity is achieved through a random process of DNA rearrangement, which involves the recombination of the germline V, D, and J gene segments and the deletion and insertion of nucleotides at the V(D)J junctions. While the theoretical diversity of different TCRs is estimated to be as high as [2], the realized diversity in an individual is much smaller, typically ranging between and [3].
At the molecular level, TCRs interact with peptides presented on the major histocompatibility complex (MHC), a complex commonly referred to as pMHC. Although the interaction between pMHC and TCR is highly specific, a single TCR can often recognize multiple pMHC complexes. Indeed, some TCRs have been shown to recognize up to a million different epitopes [4]. This multivalency is necessary to ensure that the realized diversity in one individual can recognize a significantly broader array of potential antigens.
2 T cell receptor specificity prediction.
The precise prediction of TCR-pMHC binding is essential for accurately quantifying and predicting immune responses. If effectively represented, it has the potential to transform the field of personalized medicine. For instance, the accurate identification of epitopes recognized by expanded TCR clones can aid in identifying the auto-antigens involved in T-cell-associated autoimmune diseases [5], assessing responses to vaccines or identifying the pathogenic agents responsible for eliciting T-cell responses [6]. In the context of cancer, an improved predictive power of TCR specificity can not only aid in the design of more effective T cell-based therapies [7], but also minimize toxic side-effects produced by TCRs off-target binding [8].
However, as experimental methods cannot encompass the vast space of potential TCRs and epitopes, significant emphasis has been placed on the development of reliable computational methods to predict TCR specificity. Existing methods can accurately classify in-distribution samples, i.e. they can predict TCR binding to epitopes already encountered by the model [9]. However, the pivotal challenge is to develop models with the capacity to generalize to novel epitopes. A major challenge stems from the scarcity of datasets containing experimentally validated TCR-epitope interactions, and in particular, regarding the diversity of sampled epitopes.
3 Limitations of available datasets.
TCR specificity data can be collected from various databases, such as the VDJdb [10], with over 70,000 TCR sequences and 1100 different epitopes as of December 2023, and McPas-TCR [11], with a manually curated set of 40,000 pairs. Newer datasets are also rapidly becoming available, such as the MIRA dataset, published during the COVID pandemic and including over 135,000 TCRs binding various COVID-19 epitopes [6].
However, current datasets exhibit serious limitations. First, while bulk sequencing of T cells is high-throughput and cost-effective, it cannot detect paired and chain sequences. New single-cell technologies can generate paired-chain data, yet they are costly and remain relatively underrepresented in public datasets. Currently, only a minor fraction of samples in VDJdb and none in the MIRA datasets provide paired-chain data. Furthermore, the experimental methods predominantly rely on known target pMHC complexes, skewing the datasets towards TCRs that recognize a limited number of epitopes, predominantly of viral origin and associated with the 3-6 most common HLA alleles. Finally, the datasets also show a significant bias in epitope diversity, with just 1̃00 antigens accounting for 70% of TCR-antigen pairs [12].
The lack of negative data in T cell sequencing, which focuses primarily on pMHC-labeled cells, further challenges the development of accurate supervised machine learning models. Different approaches are typically used to artificially generate non-binding TCR-epitope pairs, ranging from shuffling TCR-epitope pairs, to using naive TCR sequences, or decoy datasets. However, each of these methods introduces its own biases, and careful consideration is needed in their application to ensure the generation of negative data that accurately reflects true non-binding interactions [13, 14].
4 Evolution of TCR specificity prediction models.
Since the first release of TCR-pMHC binding data in 2017 [15], multiple studies have undertaken the challenge of modeling TCR specificity, with models broadly categorized into three groups: unsupervised clustering methods, supervised classifiers, and protein language models. Namely, in the early years of TCR-pMHC modeling (2017-2019), when the scarcity of labeled data posed challenges for training supervised models, simple clustering algorithms demonstrated the feasibility of predicting TCR specificity from sequences. In 2020 and 2021, with the increased availability of data, there was a surge in supervised models ranging from simple classifiers to deep neural networks. In the last couple of years, the breakthrough of Large Language Models, such as OpenAI’s generative pre-trained transformer (GPT) models [16], has facilitated the emergence of Protein Language Models (PLMs) based on similar principles [17, 18, 19, 20]. The number of PLMs is rapidly increasing, with some of them being specifically trained on TCR sequences [21, 22].
Fig. 1 illustrates the evolution of modeling approaches, prompting us to refer to the different waves as generations. In what follows, we describe these three model generations and summarize the main characteristics of prominent representatives of each.
4.1 1st generation: unsupervised clustering
The initial efforts in TCR specificity prediction employed unsupervised clustering methods, under the assumption that sequence similarity, or more precisely, similarity of sequence features, correlates with specificity similarity. Under this hypothesis, clusters of TCRs with similar sequences are expected to bind to the same targets. Typically, these approaches first established a distance measure, and applied the K-nearest neighbor method to determine labels for test samples based on the closest training set samples.
A simple proof of concept was provided by De Neuter et al [23]. On a dataset comprising only two HIV epitopes, a random tree classifier was trained using TCR features including V and J segments, CDR3 sequence length, mass, amino acid counts, as well as biochemical features, such as CDR3 basicity, hydrophobicity, helicity, and isoelectric point. While the applications of this model were limited, it elegantly showed that specificity prediction from sequences is feasible. An advanced version of the model was published in 2019 as TCRex [24].
TCRdist, the most established TCR distance measure, was introduced by Dash et al. in 2017 [15]. It is based on the Hamming distance–the number of edits required to transform one sequence into another–enhanced with a gap penalty, and it assigns greater weight to the CDR loops due to their importance for epitope binding. Using this distance measure, it was shown that TCRs with similar specificity cluster together, and the specificity of uncharacterized TCRs can be correctly predicted based on their proximity to training sequences. A newer version of this distance measure incorporating several new features has been introduced as TCRdist3 [25].
In the same year, Glanville and colleagues released GLIPH (Grouping of Lymphocyte Interactions by Paratope Hotspots) [26], a TCR clustering model based on both global similarity and local motif similarity. The model establishes an edge between TCRs that either differ by fewer than two amino acids or share a sequence motif enriched more than 10-fold relative to naive repertoires. Clusters are then determined as communities in the resulting graph.
The success of these early models demonstrated that the TCR sequence encodes specificity information and led to the identification of a set of features, i.e. high-level sequence descriptors, edit distance, or motif sharing, which have since been utilized by subsequent, more complex models. Interestingly, even with the increased availability of data, distance-based approaches, e.g. TCRMatch [27], GIANA [28], iSMART [29], ELATE [30], etc, have continued to be developed and demonstrate competitive performance on many tasks. However, while clustering approaches are straightforward and effective in environments with limited data, they struggle to accurately represent complex nonlinear interactions and outliers. As more data became increasingly accessible through public databases such as VDJdb, these approaches soon began to be replaced by more sophisticated supervised methods.
4.2 2nd generation: supervised models
In the domain of supervised models, clear distinctions emerge in both the modeling approach and the formulation of the prediction task. The modeling aspect spans a range from various nonparametric machine learning models to neural network architectures. As for the prediction task, there are two primary methods. The first treats known epitopes as distinct classes, to which TCRs are assigned. This method, used by all unsupervised clustering methods and approximately one-third of supervised models, does not utilize epitope information as input, and hence, it cannot generalize to unknown epitopes. Conversely, models that explicitly incorporate epitopes as input attempt to predict the binding probability between any given TCR and epitope.
Early models predominantly employed non-parametric methods and treated epitopes as class labels for TCRs. Notable examples include TCRGP [31] and SETE (Sequence-based Ensemble learning approach for TCR Epitope binding prediction) [32]. TCRGP searches for similarities between TCRs using a Gaussian Process classifier with a squared exponential kernel function based on a BLOSUM encoding, i.e. an amino acid-based encoding extracted from the Blocks Substitution Matrix [33]. SETE utilizes k-Mers of the CDR3 sequence as input features and adopts an ensemble learning approach based on decision trees to classify TCR sequences into epitope-binding classes.
The bulk of supervised models, however, employs neural network architectures. One of the first attempts, NetTCR [34], employed convolutional layers and separated input streams for TCR and epitope sequences, using BLOSUM encodings to vectorize amino acid sequences. Due to the low number of available training sequences at the time, the model performance was moderate, although improved versions with enhanced accuracy have been released afterwards [35, 36].
TcellMatch [37], another early neural network model, featured a variety of layer types, including self-attention, Gated Recurrent Units (GRUs), Long Short-Term Memory (LSTM), and convolutional layers. The authors explored the use of paired chain data versus only TCR data, and experimented with adding covariate data such as transcriptome and surface protein expression from single-cell experiments. Furthermore, the explicit encoding of epitopes was compared with using epitopes as class labels, with the latter method showing an improved performance.
Published shortly afterward, ImRex [13] introduced the novel concept of representing TCRs and epitopes as visual interaction maps, facilitating the use of established computer vision techniques. ImRex was one of the first models capable of making predictions for unseen epitopes, though its performance was somewhat limited.
Both DeepTCR [38] and pMTnet [39] exploited autoencoders to generate meaningful representations of TCR sequences in a latent space, later used as input for a classifier. This has the advantage of permitting the use of additional unlabeled data to train the encoder and decoder. DeepTCR employed variational autoencoders with convolutional layers, and used the TCR CDR3 sequence along with V and J gene information. The key difference in pMTnet is the use of stacked autoencoders instead of variational ones, and the explicit encoding of the pMHC molecule through a re-implemented version of the netMHCpan [40], an MHC-I binding machine-learning model.
Published in 2021, TITAN [41] is a supervised neural network inspired by drug sensitivity prediction models [42, 43]. The architecture employs convolutional layers, self-attention, and multi-head context-attention layers. Importantly, TITAN experimented with encoding peptides with SMILES (Simplified Molecular Input Line-Entry System), a linear and readable format to represent molecules atom-wise, which enables an efficient token-based input of atomic units into a neural network. Encoding epitopes as SMILE strings resulted in two significant benefits. First, due to the multiple paths for traversing a molecular graph, the same molecule can be encoded in various equivalent ways as a SMILES string, a property that was leveraged in TITAN as an effective data augmentation strategy. Second, Second, it enabled the pretraining of the network using protein-compound interactions, resulting in a substantial enhancement of the model’s performance. This is an example of transfer learning, where large amounts of related data are used to improve predictions in scenarios with limited data availability.
The key insights gained from these models are that a good latent sequence representation can significantly improve a model’s performance. Additionally, transfer learning further contributes to boosting the accuracy of models. The pretrained PLMs developed in the 3rd generation and discussed in Section 4.3 further capitalize on these principles.
4.3 3rd generation: Protein language models.
Recently, attention-based Transformer models have garnered substantial public interest due to the release of language models fine-tuned for conversational tasks, such as ChatGPT, Bard, Perplexity, and others. Although these AI-based chatbots have brought the field into the public eye, the use of self-supervised pre-training on unlabeled data has been driving advancements in natural language processing (NLP) for years.
Attention-based Transformers were first introduced in 2017 [44]. It soon became evident that these architectures could be trained on millions of lines of unlabeled text using proxy tasks such as masked word prediction or next word prediction [45, 46]. Following the pre-training phase, the models are typically fine-tuned with labeled datasets for specific text-based downstream tasks.
One of the key concepts contributing to the success of these models was the shift from one-hot encodings, a transformation that simply converts text into vectors without preserving any semantic or contextual word information, to context-informed word representations that capture the meaning and similarity of words based on their contextual usage. Examples of these are Word2Vec [47], and more recently, Transformer architectures such as GPT (Generative pre-trained Transformer) [46, 48, 49], BERT (Bidirectional Encoder Representations from Transformers) [45], Transformer-XL [50], and XLNet [51]. Transformer architectures, when trained on extensive text corpora, encounter words in diverse contexts, which enables them to construct highly informative latent word representations of significant relevance for a wide range of downstream tasks.
Learning the language of biology: With the success of Transformer architectures in NLP tasks, they were quickly adapted to various biological tasks, such as biomedical text mining [52] and genomic sequence analysis [53]. In the context of protein modeling, In protein modeling, the sequential order of amino acids in proteins follows rules determined by chemical properties, such as polarity, charge, and hydrophilicity. This is analogous to how grammatical rules govern the arrangement of words in sentences. Indeed, since the year 2020, a variety of Transformer models have emerged that, trained on extensive protein sequence datasets, are demonstrating competitive performance in various protein-related tasks, such as predicting protein homology, structure, function, and interactions [54, 55, 56, 57, 58, 59, 60].
4.3.1 Language models for TCR specificity prediction
With PLMs becoming increasingly popular, the first attempts to apply them to the TCR specificity prediction problem were soon made. For example, TCR-BERT [21] is a BERT-based model that was pre-trained on a dataset of 88,403 TCR and TCR sequences using a masked token prediction task. The weights were fine-tuned on an antigen classification task, and the resulting TCR embeddings were reduced to 50 dimensions using principal component analysis (PCA), followed by classification with a support vector machine (SVM).
More recently, STAPLER [22], another BERT-based architecture, was trained on an even larger unlabeled dataset of almost 80M random pairs of TCR, TCR, and peptide sequences using similarly a masked token prediction task.
Although not a transformer model, CatELMo [61] exploits ELMo (Embeddings from Language Models) [62], a bi-directional context-aware LSTM-based language model, for TCR modeling. CatELMo is trained on more than four million TCR sequences collected from ImmunoSEQ [6], and achieves improved performance compared to traditional TCR and epitope sequence embeddings, such as BLOSUM.
All these models leverage transfer learning by utilizing unlabeled TCR sequences for pre-training. Alternatively, transfer learning can also be utilized by leveraging existing large PLMs that have been trained on vast collections of unlabeled protein sequences. For instance, TCRconv [63] employs ProtBERT [56] embeddings as input and processes them through a convolutional and a linear layer for the downstream task of epitope classification. When considering this approach, a relevant question is whether is better to use a large, general PLM or a domain-specific PLM. A recent study showed that embeddings from ESM2 [60], a general PLM (Fig. 2), and TCR-BERT [21] yielded comparable results in a TCR classification task [64]. This study further showed that general models trained on broader collections of heterogeneous protein sequences can outperform their domain-specific counterparts in data-rich environments. This finding challenges the prevailing assumption that domain-specific models offer superior performance in specialized tasks. Indeed, as each PLM encodes information differently, the selection of PLM embeddings and the design of the coupled architecture have to be carefully optimized for each downstream application.
5 Outlook
The accurate prediction of TCR specificity is crucial for multiple clinical applications such as the development of safer and more efficient immunotherapies. It can also deepen our understanding of autoimmune diseases. Despite the progress, current TCR prediction models face significant challenges due to insufficient and often biased data, notably concerning epitope information. This limitation restricts the models’ capacity to generalize to new epitopes, which is critical, for instance, to predict potential cross-reactive events in T cell-based therapies.
In recent years, we have seen the rapid release and evolution of machine learning models exploiting supervised and unsupervised learning approaches. The advent of protein language models (PLMs), which leverage transfer learning to enhance performance in data-scarce domains, is revolutionizing the field and showing breakthrough performances in various protein-related tasks. However, substantial work is still required before these models can reliably perform highly specific tasks, such as predicting TCR specificity to novel epitopes. Importantly, alongside the rapid development of new models, targeted efforts to understand the strengths and weaknesses of each model need to be undertaken. In this direction, the ImmRep 2022 Workshop [9] has initiated efforts to create a common benchmark dataset and to conduct yearly competitions to facilitate rigorous model comparisons–a critical advancement for the field.
The interpretation of existing models is also necessary to improve our understanding of TCR-epitope interactions. Many machine learning models, especially PLMs, are inherently non-interpretable due to the encoding of amino acid information in highly abstract and complex latent spaces. The challenge of deriving meaning and explanations from black-box models presents a crucial challenge that could be addressed by leveraging recent advancements in Interpretable Machine Learning. For instance, the Automated Concept-based Explanation (ACE) [65] aims to automatically identify key concepts, i.e. high-level, human-understandable features, that influence model decisions. Applied to clusters of samples in the latent space, it could be used to explore the biochemical properties that govern cluster assignments.
Attention mechanisms can also be leveraged to capture structurally important residue pairs that contribute to TCR-epitope binding [66, 67] or predict protein structural properties [68]. For instance, a recent study categorized residues into groups with high and low attention values. The analysis revealed that residues receiving more attention often had distinct structural properties, and they were statistically more likely to form hydrogen bonds within the CDR3 region [68]. Alternatively, black-box models can be employed to generate plausible hypotheses that are later investigated using post hoc interpretable methods. An example of this approach is DECODE [69], a user-centric interpretable pipeline to extract human-comprehensible rules from any black-box TCR predictive model. DECODE exploits Anchors, a model-agnostic approach to approximate the decision boundary of any machine learning model and identify local, sufficient conditions for predictions [70]. However it is addressed, interpreting model predictions is not just a technical necessity but a fundamental requirement for clinical trust and application [71, 72, 73].
Other open questions demand nuanced investigation, such as understanding how PLMs organize information in the latent space. Counterintuitively, recent research has shown that the selecting of an appropriate PLM embedding layer for downstream applications necessitates optimization, as in some instances, earlier layers demonstrate superior representation capabilities for specific tasks [64]. Gaining a deeper insight into what and how information is encoded could enable more focused and efficient development and deployment of models for particular tasks.
Finally, many computational methods designed for TCR specificity prediction can be adapted to model B cell receptor (BCR) antigen binding. Indeed, both statistical [74], machine learning [75], and transformer-based approaches [76] have been developed to jointly model TCR and BCR repertoires. While not the primary focus of this paper, BCR specificity prediction faces similar data and computational challenges as TCR specificity prediction. A particularly interesting question is whether PLM-based models can effectively capture both TCR and BCR characteristics, or if specialized models might yield higher accuracy.
In summary, the field of computational immunology, particularly regarding TCR and BCR specificity prediction, is rapidly evolving, yet key questions and challenges persist. Addressing these is crucial not only for deepening our understanding of the immune system but also for paving the way for groundbreaking clinical applications.
6 Funding and declarations
The authors acknowledge funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie Actions (813545) and the ICT-2018-2 Program (826121). Additional support was received from the Swiss National Science Foundation through the Sinergia program (CRSII5 193832) and Project Funding (192128).
During the preparation of this work, the authors used ChatGPT to improve language clarity and readability. After using this service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
7 Appendix
Survey of TCR specificity prediction models.
| Epitope | |||
| Model | Year | Type | explicitly |
| encoded | |||
| TCRdist [15] | 2017 | Unsupervised | No |
| Gliph [26] | 2017 | Unsupervised | No |
| DeNeuter [23] | 2018 | Unsupervised | No |
| NetTCR [77] | 2018 | Supervised | Yes |
| TCRex [24] | 2019 | Unsupervised | No |
| SETE [78] | 2020 | Supervised | No |
| TcellMatch [79] | 2020 | Supervised | Yes |
| ImRex [13] | 2020 | Supervised | Yes |
| iSMART [29] | 2020 | Unsupervised | No |
| ERGO [80] | 2020 | Supervised | Yes |
| TCRdist3 [25] | 2021 | Unsupervised | No |
| TCRGP [81] | 2021 | Supervised | No |
| ERGO2 [82] | 2021 | Supervised | Yes |
| TCRMatch [27] | 2021 | Unsupervised | No |
| pMTnet [39] | 2021 | Supervised | Yes |
| TITAN [41] | 2021 | Supervised | Yes |
| NetTCR-2.0 [35] | 2021 | Supervised | Yes |
| TCRAI [83] | 2021 | Supervised | No |
| DeepTCR [38] | 2021 | Supervised | No |
| TCR-BERT [21] | 2021 | Language | No |
| GIANA [28] | 2021 | Unsupervised | No |
| ELATE [30] | 2021 | Unsupervised | No |
| DLpTCR [84] | 2021 | Supervised | Yes |
| ATM-TCR [85] | 2022 | Supervised | Yes |
| TCRconv [63] | 2022 | Language | No |
| diffRBM [86] | 2023 | Supervised | No |
| PanPep [87] | 2023 | Supervised | Yes |
| catELMo [61] | 2023 | Language | Yes |
| STAPLER [22] | 2023 | Language | Yes |
| NetTCR-2.2 [36] | 2023 | Supervised | Yes |
| TCR-H [88] | 2023 | Supervised | Yes |
| epiTCR [89] | 2023 | Supervised | Yes |
| GGNpTCR [90] | 2023 | Supervised | Yes |
| Rehman Khan [91] | 2023 | Language | No |
| TAPIR [92] | 2023 | Supervised | Yes |
| BERTrand [93] | 2023 | Language | Yes |
| MITNet [94] | 2023 | Supervised | No |
| SC-AIR-BERT [76] | 2023 | Language | No |
| MixTCRpred [95] | 2023 | Language | Yes |
| TSPred [66] | 2023 | Supervised | Yes |
| Koyama et al [68] | 2023 | Language | Yes |
| EPIC-TRACE [96] | 2023 | Language | Yes |
Both an epitope classification model and a model that encodes the epitope sequences are presented.
diffRBM trains two different models: one for immunogenicity prediction, which explicitly models the epitope sequence, and a second for TCR epitope specificity prediction, which does not consider the epitope sequence.
The basic model is epitope-specific; however, a pan-epitope model that encodes the sequence of the epitope is also presented.